# SKADA-Bench: Benchmarking Unsupervised Domain Adaptation Methods with Realistic Validation

### v1

@@ D.1 Results per datasets
ClassRegOT 0.66 $\pm$ 0.09 0.44 $\pm$ 0.07 0.66 $\pm$ 0.07 0.4 $\pm$ 0.06 0.36 $\pm$ 0.03 0.32 $\pm$ 0.06 0.62 $\pm$ 0.08 0.69 $\pm$ 0.08 0.53 $\pm$ 0.08 0.52 $\pm$ 0.07 12.78
LinOT 0.76 $\pm$ 0.1 0.55 $\pm$ 0.06 0.79 $\pm$ 0.04 0.48 $\pm$ 0.09 0.46 $\pm$ 0.06 0.46 $\pm$ 0.07 0.71 $\pm$ 0.04 0.76 $\pm$ 0.09 0.53 $\pm$ 0.07 0.61 $\pm$ 0.07 3.78
MMD-LS 0.68 $\pm$ 0.11 0.51 $\pm$ 0.05 0.72 $\pm$ 0.05 0.49 $\pm$ 0.06 0.39 $\pm$ 0.04 0.41 $\pm$ 0.06 0.62 $\pm$ 0.07 0.69 $\pm$ 0.08 0.54 $\pm$ 0.1 0.56 $\pm$ 0.07 7.67
Subspace JPCA 0.62 $\pm$ 0.04 0.48 $\pm$ 0.09 0.71 $\pm$ 0.03 0.51 $\pm$ 0.04 0.36 $\pm$ 0.05 0.39 $\pm$ 0.08 0.63 $\pm$ 0.06 0.64 $\pm$ 0.05 0.56 $\pm$ 0.09 0.54 $\pm$ 0.06 5.44
SA 0.65 $\pm$ 0.07 0.59 $\pm$ 0.05 0.73 $\pm$ 0.11 0.39 $\pm$ 0.09 0.32 $\pm$ 0.07 0.28 $\pm$ 0.05 0.59 $\pm$ 0.05 0.69 $\pm$ 0.06 0.5 $\pm$ 0.04 0.53 $\pm$ 0.07 11.06
TCA 0.3 $\pm$ 0.1 0.26 $\pm$ 0.06 0.25 $\pm$ 0.04 0.26 $\pm$ 0.03 0.22 $\pm$ 0.04 0.23 $\pm$ 0.06 0.25 $\pm$ 0.05 0.36 $\pm$ 0.06 0.26 $\pm$ 0.06 0.27 $\pm$ 0.06 19.56
TSL 0.26 $\pm$ 0.02 0.25 $\pm$ 0.04 0.25 $\pm$ 0.01 0.24 $\pm$ 0.04 0.22 $\pm$ 0.07 0.26 $\pm$ 0.02 0.27 $\pm$ 0.02 0.25 $\pm$ 0.05 0.25 $\pm$ 0.05 0.25 $\pm$ 0.03 20.11
Other JDOT 0.59 $\pm$ 0.08 0.37 $\pm$ 0.06 0.59 $\pm$ 0.07 0.39 $\pm$ 0.05 0.28 $\pm$ 0.05 0.3 $\pm$ 0.04 0.54 $\pm$ 0.09 0.62 $\pm$ 0.05 0.51 $\pm$ 0.08 0.46 $\pm$ 0.06 15.56
OTLabelProp 0.63 $\pm$ 0.06 0.41 $\pm$ 0.05 0.64 $\pm$ 0.05 0.41 $\pm$ 0.04 0.35 $\pm$ 0.07 0.33 $\pm$ 0.06 0.59 $\pm$ 0.06 0.66 $\pm$ 0.04 0.5 $\pm$ 0.09 0.5 $\pm$ 0.06 13.11
D.2 Impact of the base estimators on the simulated datasets
As mentioned in the main paper, it is possible to partly compensate for the shift by choosing the right base estimator. In this part, we provide the results on the Simulated dataset for three different base estimators: Logistic Regression (LR) in Table D.9, SVM in Table D.10, and XGBoost in Table D.11. Observing the two first rows for covariate shift, we see that with LR (Table D.9), there is a significant drop in performance between training on the source v.s. training on the target ($\sim 10\%$), while using SVC (Table D.10) only leads to a drop ($\sim 3\%$). Finally, using XGBoost (Table D.11) maintains the performance. The reweighting DA methods help compensate for the shift when using a simpler LR estimator. However when using an SVC, as shown in the main paper, the reweighting does not help to compensate for the covariate shift. If we look at the other shifts, the problem is harder. The subspace methods help with subspace shift, and the mapping methods help with the conditional shift.
@@ D.2 Impact of the base estimators on the simulated datasets
Table D.9: Accuracy score for simulated datasets compared for all the methods with LR. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Table D.10: Accuracy score for simulated datasets compared for all the methods with SVC. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Table D.11: Accuracy score for simulated datasets compared for all the methods with XGBoost. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
D.3 Unrealistic validation with supervised scorer
Table D.12 shows the results when we choose the supervised scorer that is when validating on target labels. It is important to highlight that this choice is impossible in real life applications due to the lack of target labels.
When using the target labels, the method’s parameters are better validated. This can be seen by the significant increase in the table (blue values), which are numerous in this table compared to the one with the selected realistic scorer. For example, the method MMDTarS, which is made for Target shift, compensates all the shift simulated covariate shifts when we select the model with a supervised scorer.
When looking at the rank, 11 DA methods have a higher rank than Train Src compared to 8 when using realistic scorer.
Table D.12: Accuracy score for all datasets compared for all the methods for simulated and real-life datasets. In this table, each DA method is validated with the supervised scorer. The color indicates the amount of the improvement. A white color means the method is not statistically different from Train Src (Train on source). Blue indicates that the performance improved with the DA methods, while red indicates a decrease. The darker the color, the more significant the change.
@@ D.3 Unrealistic validation with supervised scorer
Train Src 0.88 0.85 0.66 0.19 0.59 0.56 0.54 0.59 0.7 0.72 0.91 0.55 11.27
Train Tgt 0.92 0.93 0.82 0.98 0.88 0.8 0.96 1.0 0.73 1.0 0.97 0.64 1.12
Reweighting Dens. RW 0.89 0.87 0.67 0.2 0.59 0.56 0.54 0.59 0.7 0.76 0.91 0.55 11.05
Disc. RW 0.86 0.84 0.73 0.23 0.58 0.53 0.54 0.62 0.69 0.78 0.91 0.56 12.19
Gauss. RW 0.89 0.86 0.65 0.21 0.2 0.44 0.11 0.54 0.6 0.51 0.46 0.25 20.03
KLIEP 0.89 0.88 0.66 0.2 0.59 0.56 0.54 0.58 0.7 0.75 0.92 0.55 10.93
KMM 0.9 0.87 0.67 0.2 0.58 0.55 0.53 0.71 0.66 0.75 0.92 0.54 12.12
NN RW 0.89 0.86 0.67 0.15 0.58 0.55 0.55 0.59 0.66 0.72 0.91 0.54 13.79
MMDTarS 0.88 0.93 0.66 0.27 0.59 0.56 0.52 0.59 0.7 0.74 0.91 0.55 10.99
Mapping CORAL 0.66 0.84 0.66 0.19 0.6 0.57 0.62 0.75 0.7 0.72 0.92 0.62 8.34
MapOT 0.87 0.63 0.81 0.14 0.54 0.51 0.6 0.77 0.67 0.63 0.84 0.47 14.87
EntOT 0.89 0.61 0.82 0.47 0.61 0.58 0.63 0.88 0.68 0.81 0.87 0.53 9.22
ClassRegOT 0.91 0.59 0.82 0.15 0.61 0.59 0.66 0.98 0.68 0.89 0.9 0.52 6.40
LinOT 0.89 0.81 0.81 0.19 0.6 0.58 0.65 0.88 0.71 0.81 0.91 0.6 5.68
MMD-LS 0.64 0.79 0.81 0.77 0.59 0.56 0.56 0.97 0.68 0.86 0.88 0.58 9.09
Subspace JPCA 0.88 0.85 0.66 0.19 0.59 0.56 0.56 0.83 0.7 0.8 0.9 0.55 9.13
SA 0.74 0.81 0.8 0.13 0.6 0.58 0.56 0.93 0.7 0.91 0.89 0.59 7.26
TCA 0.44 0.47 0.53 0.5 0.04 NA 0.11 0.56 0.6 0.43 NA 0.27 20.29
TSL 0.88 0.85 0.66 0.86 0.59 0.21 0.3 0.7 0.7 0.61 0.88 0.26 15.12
Other JDOT 0.72 0.57 0.82 0.14 0.61 0.51 0.64 0.77 0.68 0.63 0.8 0.46 13.43
OTLabelProp 0.9 0.76 0.81 0.14 0.61 0.56 0.64 0.89 0.67 0.69 0.86 0.51 11.14
DASVM 0.89 0.86 0.65 0.14 NA NA NA 0.68 NA 0.78 0.88 NA 12.39
D.4 Comparions between supervised and unsupervised scorers

### v2

@@ E.1 Results per datasets
ClassRegOT 0.61 $\pm$ 0.07 0.43 $\pm$ 0.07 0.71 $\pm$ 0.07 0.4 $\pm$ 0.05 0.35 $\pm$ 0.04 0.31 $\pm$ 0.07 0.62 $\pm$ 0.06 0.7 $\pm$ 0.05 0.49 $\pm$ 0.06 0.51 $\pm$ 0.06 12.33
LinOT 0.76 $\pm$ 0.1 0.55 $\pm$ 0.05 0.79 $\pm$ 0.04 0.48 $\pm$ 0.09 0.46 $\pm$ 0.06 0.46 $\pm$ 0.07 0.71 $\pm$ 0.04 0.76 $\pm$ 0.09 0.53 $\pm$ 0.07 0.61 $\pm$ 0.07 3.78
MMD-LS 0.66 $\pm$ 0.1 0.46 $\pm$ 0.13 0.63 $\pm$ 0.16 0.35 $\pm$ 0.11 0.39 $\pm$ 0.05 0.34 $\pm$ 0.1 0.65 $\pm$ 0.08 0.66 $\pm$ 0.12 0.49 $\pm$ 0.08 0.51 $\pm$ 0.1 10.11
Subspace JPCA 0.57 $\pm$ 0.07 0.39 $\pm$ 0.05 0.72 $\pm$ 0.06 0.47 $\pm$ 0.07 0.31 $\pm$ 0.05 0.33 $\pm$ 0.03 0.64 $\pm$ 0.06 0.68 $\pm$ 0.04 0.49 $\pm$ 0.09 0.51 $\pm$ 0.06 12.78
SA 0.74 $\pm$ 0.1 0.58 $\pm$ 0.09 0.8 $\pm$ 0.03 0.48 $\pm$ 0.07 0.39 $\pm$ 0.04 0.4 $\pm$ 0.06 0.66 $\pm$ 0.1 0.73 $\pm$ 0.07 0.54 $\pm$ 0.06 0.59 $\pm$ 0.07 6.33
TCA 0.31 $\pm$ 0.09 0.26 $\pm$ 0.03 0.24 $\pm$ 0.07 0.24 $\pm$ 0.1 0.27 $\pm$ 0.05 0.26 $\pm$ 0.07 0.27 $\pm$ 0.06 0.31 $\pm$ 0.1 0.29 $\pm$ 0.1 0.27 $\pm$ 0.07 19.50
TSL 0.24 $\pm$ 0.04 0.26 $\pm$ 0.05 0.27 $\pm$ 0.05 0.26 $\pm$ 0.03 0.23 $\pm$ 0.07 0.27 $\pm$ 0.04 0.26 $\pm$ 0.01 0.24 $\pm$ 0.09 0.22 $\pm$ 0.03 0.25 $\pm$ 0.05 20.00
Other JDOT 0.57 $\pm$ 0.07 0.37 $\pm$ 0.06 0.61 $\pm$ 0.05 0.39 $\pm$ 0.07 0.28 $\pm$ 0.04 0.3 $\pm$ 0.03 0.53 $\pm$ 0.09 0.62 $\pm$ 0.05 0.49 $\pm$ 0.05 0.46 $\pm$ 0.06 17.00
OTLabelProp 0.63 $\pm$ 0.06 0.4 $\pm$ 0.06 0.64 $\pm$ 0.05 0.41 $\pm$ 0.04 0.35 $\pm$ 0.07 0.33 $\pm$ 0.06 0.59 $\pm$ 0.06 0.66 $\pm$ 0.04 0.5 $\pm$ 0.09 0.5 $\pm$ 0.06 13.11
Table 16: Accuracy score for BCI dataset for each shift compared for all the deep DA methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
1 2 3 4 5 6 7 8 9 Mean Rank
Train Src 0.59 0.29 0.67 0.43 0.38 0.29 0.67 0.76 0.67 0.53 2.61
Train Tgt 0.57 0.43 0.71 0.53 0.34 0.29 0.62 0.69 0.79 0.55 2.17
DANN 0.52 0.24 0.55 0.38 0.4 0.22 0.38 0.66 0.64 0.44 4.44
DeepCORAL 0.55 0.26 0.66 0.36 0.48 0.33 0.57 0.72 0.66 0.51 2.72
DeepJDOT 0.53 0.34 0.64 0.43 0.43 0.43 0.57 0.62 0.6 0.51 3.17
E.2 Impact of the base estimators on the simulated datasets
As mentioned in the main paper, it is possible to partly compensate for the shift by choosing the right base estimator. In this part, we provide the results on the Simulated dataset for three different base estimators: Logistic Regression (LR) in Table 17, SVM in Table 18, and XGBoost in Table 19. Observing the two first rows for covariate shift, we see that with LR (Table 17), there is a significant drop in performance between training on the source v.s. training on the target ($\sim 10\%$), while using SVC (Table 18) only leads to a drop ($\sim 3\%$). Finally, using XGBoost (Table 19) maintains the performance. The reweighting DA methods help compensate for the shift when using a simpler LR estimator. However when using an SVC, as shown in the main paper, the reweighting does not help to compensate for the covariate shift. If we look at the other shifts, the problem is harder. The subspace methods help with subspace shift, and the mapping methods help with the conditional shift.
@@ E.2 Impact of the base estimators on the simulated datasets
Table 17: Accuracy score for simulated datasets compared for all the methods with LR. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Table 18: Accuracy score for simulated datasets compared for all the methods with SVC. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Table 19: Accuracy score for simulated datasets compared for all the methods with XGBoost. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
E.3 Unrealistic validation with supervised scorer
Table 20 shows the results when we choose the supervised scorer that is when validating on target labels. It is important to highlight that this choice is impossible in real life applications due to the lack of target labels. When using the target labels, the method’s parameters are better validated. This can be seen by the significant increase in the table (blue values), which are numerous in this table compared to the one with the selected realistic scorer. For example, the method MMDTarS, which is made for Target shift, compensates all the shift simulated covariate shifts when we select the model with a supervised scorer. When looking at the rank, 11 DA methods have a higher rank than Train Src compared to 9 when using realistic scorer. The findings hold for Deep DA where the accuracy in Table 21 is overall better than when using unsupervised scorers.
Table 20: Accuracy score for all datasets compared for all the methods for simulated and real-life datasets. In this table, each DA method is validated with the supervised scorer. The color indicates the amount of the improvement. A white color means the method is not statistically different from Train Src (Train on source). Blue indicates that the performance improved with the DA methods, while red indicates a decrease. The darker the color, the more significant the change.
@@ E.3 Unrealistic validation with supervised scorer
Train Src 0.88 0.85 0.66 0.19 0.65 0.56 0.54 0.59 0.7 0.72 0.91 0.55 10.66
Train Tgt 0.92 0.93 0.82 0.98 0.89 0.8 0.96 1.0 0.73 1.0 0.97 0.64 1.55
Reweighting Dens. RW 0.89 0.87 0.67 0.2 0.65 0.56 0.54 0.59 0.7 0.76 0.91 0.55 12.20
Disc. RW 0.86 0.84 0.73 0.23 0.64 0.54 0.54 0.62 0.69 0.78 0.91 0.56 8.75
Gauss. RW 0.89 0.86 0.65 0.21 0.22 0.44 0.11 0.54 0.55 0.51 0.46 0.25 16.45
KLIEP 0.89 0.88 0.66 0.2 0.65 0.56 0.54 0.58 0.7 0.75 0.92 0.55 10.56
KMM 0.89 0.87 0.67 0.19 0.64 0.55 0.53 0.71 0.66 0.75 0.92 0.54 11.74
NN RW 0.89 0.86 0.67 0.15 0.65 0.55 0.55 0.59 0.66 0.72 0.91 0.54 9.15
MMDTarS 0.88 0.93 0.66 0.27 0.65 0.56 0.54 0.59 0.7 0.74 0.91 0.56 10.81
Mapping CORAL 0.74 0.84 0.82 0.19 0.66 0.57 0.62 0.75 0.7 0.72 0.92 0.62 5.08
MapOT 0.87 0.63 0.82 0.14 0.6 0.51 0.6 0.77 0.68 0.63 0.84 0.47 10.21
EntOT 0.89 0.61 0.82 0.47 0.66 0.58 0.63 0.88 0.68 0.81 0.87 0.53 9.40
ClassRegOT 0.91 0.59 0.82 0.15 NA 0.59 0.66 0.98 0.68 0.89 0.9 0.52 8.25
LinOT 0.89 0.81 0.81 0.19 0.66 0.58 0.65 0.88 0.71 0.81 0.91 0.61 4.06
MMD-LS 0.88 0.85 0.81 0.73 0.65 0.56 0.56 0.98 0.69 0.89 NA 0.58 8.22
Subspace JPCA 0.88 0.85 0.66 0.19 0.65 0.56 0.56 0.84 0.7 0.8 0.9 0.55 8.98
SA 0.74 0.81 0.8 0.13 0.66 0.57 0.56 0.93 0.7 0.91 0.89 0.59 7.80
TCA 0.4 0.46 0.5 0.58 0.04 0.02 0.11 0.49 0.61 0.46 0.49 0.27 17.58
TSL 0.88 0.85 0.66 0.86 0.62 0.48 0.45 0.7 0.69 0.57 0.9 0.26 15.09
Other JDOT 0.72 0.57 0.82 0.13 0.61 0.41 0.57 0.8 0.67 0.63 0.8 0.46 11.42
OTLabelProp 0.9 0.76 0.81 0.14 0.66 0.56 0.64 0.89 0.67 0.69 0.86 0.51 10.01
DASVM 0.89 0.86 0.64 0.12 NA NA NA 0.83 NA 0.76 0.86 NA 7.29
Table 21: Accuracy score compared for the Deep methods with the supervised scorer for a selection of real-life datasets.
MNIST/USPS Office31 OfficeHome BCI Rank
Train Src 0.85 0.77 0.58 0.54 6.19
Train Tgt 0.98 0.96 0.83 0.56 2.07
DeepCORAL 0.93 0.82 0.63 0.54 3.29
DAN 0.91 0.79 0.61 0.55 4.76
DANN 0.9 0.76 0.6 0.42 4.98
DeepJDOT 0.93 0.83 0.63 0.54 2.92
MCC 0.94 0.81 0.66 0.55 2.38
MDD 0.91 0.83 0.58 0.42 4.96
SPA 0.92 0.78 0.56 0.4 5.39
E.4 F1-score of benchmark
We provide in Table 22 a version of Table 2 where the performance measure reported is the F1-score. It is interesting to note that the dynamic of which methods work best and are the more robust is very similar to the accuracy performance which illustrate the robustness of the benchmark.
Table 22: F1 score for all datasets compared for all the shallow methods for simulated and real-life datasets. The color indicates the amount of the improvement. A white color means the method is not statistically different from Train Src (Train on source). Blue indicates that the score improved with the DA methods, while red indicates a decrease. The darker the color, the more significant the change.
Train Src 0.88 0.87 0.62 0.17 0.64 0.56 0.52 0.56 0.65 0.72 0.91 0.53 10.95
Train Tgt 0.92 0.92 0.82 0.98 0.89 0.8 0.96 1.0 0.69 1.0 0.97 0.63 1.70
Reweighting Dens. RW 0.88 0.88 0.62 0.16 0.61 0.56 0.52 0.55 0.65 0.72 0.91 0.52 IW 12.71
Disc. RW 0.85 0.86 0.7 0.16 0.63 0.53 0.48 0.56 0.63 0.76 0.91 0.54 CircV 8.39
Gauss. RW 0.89 0.88 0.61 0.18 0.15 0.4 0.03 0.43 0.49 0.35 0.29 0.1 CircV 16.53
KLIEP 0.88 0.88 0.62 0.17 0.64 0.55 0.52 0.56 0.65 0.72 0.91 0.52 IW 10.66
KMM 0.89 0.87 0.6 0.15 0.63 0.54 0.51 0.69 0.5 0.74 0.91 0.49 CircV 11.58
NN RW 0.89 0.88 0.63 0.14 0.64 0.55 0.52 0.56 0.64 0.71 0.91 0.5 CircV 8.22
MMDTarS 0.88 0.88 0.6 0.17 0.57 0.56 0.52 0.56 0.65 0.74 0.91 0.53 IW 11.02
Mapping CORAL 0.74 0.76 0.74 0.16 0.65 0.57 0.62 0.72 0.65 0.72 0.92 0.62 CircV 5.00
MapOT 0.72 0.65 0.82 0.02 0.59 0.5 0.61 0.76 0.59 0.63 0.84 0.47 PE 10.49
EntOT 0.71 0.67 0.82 0.12 0.63 0.57 0.59 0.83 0.49 0.75 0.85 0.53 CircV 10.15
ClassRegOT 0.74 0.66 0.81 0.11 NA 0.53 0.62 0.97 0.67 0.82 0.89 0.52 IW 6.49
LinOT 0.73 0.78 0.75 0.16 0.65 0.57 0.64 0.81 0.65 0.76 0.91 0.61 CircV 4.20
MMD-LS 0.77 0.77 0.75 0.54 0.64 0.56 0.54 0.97 0.6 0.85 NA 0.48 MixVal 7.58
Subspace JPCA 0.88 0.87 0.62 0.14 0.61 0.47 0.5 0.76 0.61 0.78 0.9 0.51 PE 9.55
SA 0.73 0.74 0.8 0.1 0.64 0.57 0.55 0.88 0.56 0.77 0.89 0.52 CircV 7.95
TCA 0.51 0.56 0.5 0.61 0.0 0.0 0.02 0.53 0.46 0.44 0.47 0.19 DEV 17.94
TSL 0.88 0.87 0.63 0.17 0.63 0.47 0.45 0.59 0.58 0.28 0.9 0.21 PE 15.46
Other JDOT 0.72 0.66 0.82 0.13 0.59 0.41 0.59 0.8 0.61 0.65 0.79 0.46 IW 10.74
OTLabelProp 0.72 0.67 0.8 0.07 0.65 0.54 0.62 0.86 0.58 0.64 0.86 0.49 CircV 10.80
DASVM 0.89 0.88 0.61 0.13 NA NA NA 0.87 NA 0.82 0.85 NA MixVal 7.12
E.5 Comparison of the rank of DA scorer
To provide a more detailed assessment of the scorers’ performance, we present a critical difference diagram of their rankings in Figure 4. The diagram highlights that the unrealistic supervised scorer significantly outperforms all others. Among the unsupervised scorers, CircV and IW achieve the best performance, with their rankings being very close and not statistically different according to a statistical test. Next, we observe a group comprising PE, DEV, and MixVal, where DEV and MixVal are also not statistically distinguishable. Finally, SND emerges as the worst-performing scorer in the benchmark.
To give a more detailed perspective, we present a visualization in Figure 5, showing the rank of each scorer for each DA method. In the right part of the figure, the supervised scorer (in pink) is consistently the top-ranked, as expected, across all methods. Similarly, CircV (in red) and IW (in orange) consistently outperform other scorers.
Figure 4: Critical difference diagram of average ranks for scorers, computed across shallow methods and shifts (lower ranks indicate better performance). Black lines between scorers indicate pairs that are not statistically different based on the Wilcoxon test.
Figure 5: Illustrations as spider plots for all methods of the accuracy on each dataset (left) and the scorers rankings (right). For methods with no accuracy results (NA in Table 2) we replace the value by 0. We provide both spider plot in the same Figure to allow a comparison of the scorer ranking while having the possibility to check the performance for each method.
E.6 Comparisons between supervised and unsupervised scorers